OpenStudy (anonymous):

Solve |x + 8| ≤ 5. x ≤ –13 or x ≥ –3??

OpenStudy (freckles):

\[|f(x)| \le 5\] => \[-5 < f(x)<5 \]

OpenStudy (anonymous):

^ luis is correct :)

OpenStudy (freckles):

\[-5 \le f(x) \le 5 = > -5 \le x+8 \le 5 => -13 \le x \le -3\]

OpenStudy (anonymous):

|x + 8| ≤ 5 -5 ≤x + 8 ≤ 5 -5 -8 ≤ x ≤ 5 -8 => -13 ≤ x ≤ -3

Directrix (directrix):

@ChinaGurl ---> This is an AND problem and not an OR problem. x ≥ -13 AND x≤ - 3 which can be written as above: -13 ≤ x ≤ -3